True or False.

The value of the integral "\displaystyle{\int_{1}^{2}x(x^2-1)^8 dx}" is an integer


1
Expert's answer
2021-08-23T15:42:54-0400

"\displaystyle{\int_{1}^{2}x(x^2-1)^8 dx}"


Let substitute, "u=x^2-1"


Hence, "\frac{du}{dx}=2x," "dx=\frac{1}{2x}du"


"=\frac{1}{2}\smallint{u^8du}"


Now solving, "\smallint{u^8du}"


Apply power rule


"\smallint{u^8du}=\frac{u^{n+1}}{n+1}" With "n=8"


"=\frac{u^9}{18}"


Undo the substitution. "u=x^2-1"


"=\frac{(x^2-1)^9}{18}"


Replacing x with 2 and 1


"=\frac{(2^2-1)^9}{18}-0=1093.5"


False because 1093.5 is not a whole valued number